OpenStudy (anonymous):

I have a math problem stating y=8-3x , but it is not done how it is writen what should i do?

OpenStudy (anonymous):

Add 3x to both sides of the equation: 3x+y=8-3x+3x you get 3x+y=8. Subtract 8 from both sides 3x+y-8+8-8 and you get the equation you want to work with: 3x+y-8=0

OpenStudy (anonymous):

so how would i work that out?

OpenStudy (anonymous):

Are you solving for the intercepts or are you doing a solution table?

OpenStudy (anonymous):

i am doing a function table the x axis 0, 1, 2, 3. and the y axis blank, 5, 2, -1

OpenStudy (anonymous):

You are looking for the y axis amount that corresponds in the function table with the 0?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

OK.....plug the 0 into the equation in place of the x and solve for y. That will give you the y coordinate that you are looking for.

OpenStudy (anonymous):

thanx